OpenStudy (anonymous):

if you have 87 out of 200, what is the percentage

OpenStudy (anonymous):

if you have 50 out of 100 what is the percentage?

OpenStudy (anonymous):

50% but that is out of 100 not 200

OpenStudy (a1234):

Do 87/200 = x/100.

OpenStudy (anonymous):

yes, that's absolutely right. Think about the word percent. What it literally means is "per" = out of, and "cent" = 100

OpenStudy (anonymous):

therefore if you have 50 out of a 100, that means you have 50 per (out of) cent (a hundred).

OpenStudy (anonymous):

So let's think about this. if you have 25 out of 50, what does that mean?

OpenStudy (anonymous):

it is 50% so then if I multiplied 87X2 I would then be able to determine the percentage?

OpenStudy (anonymous):

No, that is wrong. Suppose you have 87 out of a hundred. That would mean you have 87 percent, right? But you only have 87 out of two hundred. That means you have a lesser percentage out of a hundred, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

another way to think about it is, if you got 87 marks out of a 100, you get an A. But if you only got 87 out of 200, what grade would you get? Higher or lower?

OpenStudy (anonymous):

lower.

OpenStudy (anonymous):

That is exactly right. So, the way you calculate percentage is, you convert everything in terms of one hundred. So if you got 87 marks out of 200, that means you only got 43.5 marks out of a hundred. In other words, you got 43.5 percent or 43.5%

OpenStudy (anonymous):

in mathematics, you would write out an equation which sounds like this: "If I got 87 out of 200, what would I get out of 100?". Which is nothing but: \[\frac{ 87 }{ 200 } = \frac{ x }{ 100 }\]

OpenStudy (anonymous):

oh, thx now I understand

OpenStudy (anonymous):

cool. have fun!